OpenStudy (emilyjones284):

How do you simplify 9x^0 y^-8 over z^-8

OpenStudy (lgbasallote):

hint: \[\huge \frac{9x^0 y^{-8}}{z^{-8}} \implies \frac{9x^0 z^8}{y^8}\] i leave the simplifying to you does that help?

OpenStudy (emilyjones284):

How did you get the z^-8 and the y^-8 to be switched and then positive...also, I dont know what to do after that.

OpenStudy (lgbasallote):

\[\huge y^{-a} \implies \frac 1{y^a}\] so \[\huge y^{-8} \implies \frac{1}{y^8}\] then \[\huge \frac 1{z^{-a}} \implies z^a\] so \[\huge \frac{1}{z^{-8}} \implies z^8\] got it?

OpenStudy (emilyjones284):

Yeah I think so...is the answer 9x^0z^8 over y^8?

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

what do you think is the value of x^0?

OpenStudy (emilyjones284):

X?

OpenStudy (emilyjones284):

the value of x^0 is x or is it 0?

OpenStudy (anonymous):

it's 1. power to 0 is 1

OpenStudy (lgbasallote):

anything raised to 0 is 1

OpenStudy (emilyjones284):

So the answer would be 9xz^8 over y^8?

OpenStudy (lgbasallote):

x^0 = 1

OpenStudy (lgbasallote):

so x^0 will just disappear

OpenStudy (lgbasallote):

and why it's 1 is a long explanation

OpenStudy (anonymous):

\[\frac{a^b}{a^b} =a^{b-b}=a^0\]\[\frac{a^b}{a^b} = 1\]here's why it's 1 :P

OpenStudy (emilyjones284):

Ohhh so it would be 9z^8 over y^8?

OpenStudy (lgbasallote):

yup

OpenStudy (emilyjones284):

Yay thank you! Can you tell me what type of equation this is so I can look it up in my textbook?